solve and check the following equation for x.
x/6 - x/8=1

Answer by checkley71(8403)

X/6-X/8=1 OR 8X/48-6X/48=1 OR (8X-6X)/48=1 OR 2X/48=1 OR 2X=48 OR X=24
PROOF 24/6-24/8=1 OR 4-3=1 OR 1=1